OpenStudy (anonymous):

When you make the circle smaller, which number in the standard equation for a circle centered at the origin decreases?

OpenStudy (whpalmer4):

Circle centered at the origin has formula \[x^2+y^2= r^2\]where \(r\) is the radius. If the circle gets smaller, what changes?

OpenStudy (anonymous):

So the standard form for a circle centered at the origin is\[(x)^2+(y)^2 = r^2\] where r is the radius of the circle. So if you made that r^2 term smaller it would shrink the circle.

OpenStudy (anonymous):

The origin?

OpenStudy (anonymous):

or the radius??

OpenStudy (whpalmer4):

The origin is (0,0). It's just a point, how could it shrink?

OpenStudy (anonymous):

It would reduce the size of the circle

OpenStudy (whpalmer4):

If the circle shrinks, the circle gets smaller, right? And the circle is just the set of all points an equal distance (the radius) from the center.

OpenStudy (whpalmer4):

So the radius shrinks as the circle gets smaller...

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks for explaining guys(:

OpenStudy (jhannybean):